a balloo starts rising from ground from rest with an upward acceleration of 2 ms-2 . Just after 1 second, a stone is dropped from it.The time taken by stone to strike the ground is nearly:
1)0.3 secs
2)0.7 secs
3)1 secs
4)1.4 secs

Grade:11

2 Answers

bharat bajaj IIT Delhi
9 years ago
After 1 second,
The speed of stone = The speed of balloon at that instant
Distance from ground when stone is dropped = 1/2. 2 = 1 m
Speed of stone = 2.1 = 2 m/s upwardsg = 10
v = u + at
s = ut + 1/2 a t^2
-1 = 2t - 5t^2
5t^2 - 2t - 1 = 0t = 2 + v24 / 10
t = 0.7 sOption 2) is correct.Thanks
Bharat BajajIIT Delhi
askiitians faculty
Arun Kumar IIT Delhi
9 years ago
Hello
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty